# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

### v1

@@ Appendix C Comparative Experiment with Optimization-based Method
To further evaluate the success rate of the progress allocation module in our algorithm, we compare it against a widely-used optimization-based algorithm Gilpin et al. [2020] in a custom-built simulation environment. The baseline algorithm is employed as a sound and complete solution to accurately assess the feasibility of randomly generated test cases.
@@
The experiment is conducted within a bounded $10\times 10$ square 2D plane containing a circular obstacle. The underlying system dynamics are modeled using a double integrator. The agent starts from a randomly generated position and must complete the randomly generated STL tasks by reaching the target region within the specified time interval.
@@
In this experiment, the baseline algorithm is implemented using the open-source library stlpy Kurtz and Lin [2022] and has full knowledge of the environmental information and system dynamics, while our algorithm only has access to the trajectory dataset.
@@
To generate the trajectory dataset, we randomly sample start and end points in the environment and use the baseline algorithm to solve reach-avoid tasks. This process produces 200,000 collision-free trajectories that satisfy the system dynamics, which are then used to train the diffusion model.
@@
Following the training procedure from Janner et al. [2022], we train two diffusion models with different planning horizons, as described in Section B.2. The first model is trained on trajectory segments of length 16 for short trajectories, while the second model uses segments of length 32 to improve generalization to longer trajectories.
@@
We generate 200 feasible STL tasks for each template, as described in Section 5. The deterministic baseline algorithm is used to ensure the feasibility of these tasks. However, for templates 4 and 5, which involve multi-layer nesting of temporal operators, the baseline algorithm fails to find solutions within an acceptable time. In these cases, we still employ our algorithm’s progress allocation module to verify feasibility.
@@
In addition to the Execution Success Rate (SR) and Total Planning Time (T0) metrics described in Section 5, we introduce an additional evaluation metric:
@@
Type SR0(%) SR(%) Total Planning Time(s)$\downarrow$
@@
1 96.0 93.5 0.99$\pm$0.08 3.82$\pm$1.44
2 98.0 96.5 0.81$\pm$0.03 6.30$\pm$1.36
3 96.0 89.0 1.26$\pm$0.05 31.60$\pm$10.46
4 - 78.5 2.10$\pm$0.26 Timeout
5 - 83.0 2.57$\pm$0.10 Timeout
6 97.5 69.5 2.87$\pm$0.12 24.23$\pm$6.39
7 80.0 73.5 1.80$\pm$0.08 7.71$\pm$3.50
8 89.5 89.0 0.82$\pm$0.03 106.58$\pm$82.19
9 81.0 72.0 1.61$\pm$0.06 151.19$\pm$78.82
Table C.1: Result of Experiment in Custom-built Environment. SR0: Progress Allocation Success Rate; SR: Execution Success Rate;
The experimental results are summarized in Table C.1. Our algorithm achieves consistently high success rates across test cases generated from diverse task templates. Notably, the Execution Success Rate (SR) exceeds 70% in all scenarios, demonstrating the algorithm’s strong generalization capability for STL tasks.
@@
Appendix D Implementation Details of Experiments
D.1 Trajectory Generation
In this work, we employ diffusion model to generate only the state sequence of the trajectory to avoid the practical issues discussed in Ajay et al. [2022]. During testing, we use a simple PD controller to follow the state sequence to get the action sequence on-the-fly and generate the final actual execution trajectory, as it performs better than employing an inverse dynamics model used in Ajay et al. [2022] to generate actions.
D.2 Calculation of the Robustness Value
Besides the Boolean semantics, STL also incorporates a concept of robustness, which refers to its quantitative semantics used to measure the degree to which a signal satisfies or violates a formula. Positive robustness values signify satisfaction, while negative values indicate violation. The quantitative semantics of a STL formula $\varphi$ with respect to a signal $\mathbf{s}_{t}$ starting at time $t$ are defined as follows:
$\rho\left(\mathbf{s}_{t},\top\right)$ $=\rho_{\max},\text{ where }\rho_{\max}>0,$ (D.1)
$\rho\left(\mathbf{s}_{t},\mu\right)$ $=h_{\mu}(\mathbf{x}_{t}),$
$\rho\left(\mathbf{s}_{t},\neg\mu\right)$ $=-h_{\mu}(\mathbf{x}_{t}),$
In the experiments, we utilized the commonly adopted open-source library stlpy Kurtz and Lin [2022] to compute the robustness values of both the planned trajectories and the actual execution trajectories. In the Maze2D environment, transitions between states require relatively long trajectories, which result in extended time intervals for the corresponding STL tasks. This creates computational overhead for calculating robustness values using stlpy, exceeding the capacity of the experimental hardware.
To address this issue, we introduced an additional sampling factor, $\eta$, which represents the mapping relationship between the time length in the STL task and the actual system trajectory. Specifically, we assume that each time step in the STL task corresponds to a trajectory length of $\eta$ in the system. When computing robustness values, we sample one state from the system trajectory for every $\eta$ states, resulting in a sampled trajectory. The robustness value is then calculated based on this sampled trajectory. While this operation may introduce minor deviations in the robustness value calculations, it does not affect the validity of the experimental conclusions.
D.3 Experimental Parameter Settings
@@
Some of the parameters involved in the experiments are listed below, and their specific values are shown in Table D.1:

### v2

@@ F.5 Details of Comparative Experiment with Optimization-based Method
To further evaluate the success rate of the progress allocation module in our algorithm, we compare it against a widely-used optimization-based algorithm [4] in a custom-built simulation environment. The baseline algorithm is employed as a sound and complete solution to accurately assess the feasibility of randomly generated test cases.
@@
Experimental Settings. The experiment is conducted within a bounded $10\times 10$ square 2D plane containing a circular obstacle. The underlying system dynamics are modeled using a double integrator. The agent starts from a randomly generated position and must complete the randomly generated STL tasks by reaching the target region within the specified time interval.
@@
In this experiment, the baseline algorithm is implemented using the open-source library stlpy [2] and has full knowledge of the environmental information and system dynamics, while our algorithm only has access to the trajectory dataset.
@@
To generate the trajectory dataset, we randomly sample start and end points in the environment and use the baseline algorithm to solve reach-avoid tasks. This process produces 200,000 collision-free trajectories that satisfy the system dynamics, which are then used to train the diffusion model and the Time Predictor.
@@
To further enhance the trajectory generator’s ability to produce trajectories of varying lengths, we train two diffusion models with different horizons: one dedicated to generating shorter trajectories and the other specialized for longer trajectories. The first model is trained on trajectory segments of length 16 for shorter trajectories, while the second model uses segments of length 32 to improve generalization to longer trajectories.
@@
We generate 200 feasible STL tasks for each template, as described in Section F.2. The deterministic baseline algorithm is used to ensure the feasibility of these tasks. However, for templates 4 and 5, which involve multi-layer nesting of temporal operators, the baseline algorithm fails to find solutions within an acceptable time. In these cases, we still employ our algorithm’s progress allocation module to verify feasibility.
@@
Evaluation Metrics. In addition to the Execution Success Rate (SR) and Total Planning Time (T0) metrics described in Section 5, we introduce an additional evaluation metric:
@@
Table F.4: Result of Experiment in Custom-built Environment. SR0: Progress Allocation Success Rate; SR: Execution Success Rate;
[…]
@@
[…]
Analysis. The experimental results are summarized in Table F.4. Since the optimization-based baseline is used as an expert solver to certify the feasibility of randomly generated STL tasks, its execution success rate (SR) is naturally 100% and thus omitted from Table F.4. Our algorithm achieves consistently high success rates across test cases generated from diverse task templates. Notably, the Execution Success Rate (SR) exceeds 69% in all scenarios, demonstrating the algorithm’s strong generalization capability for STL tasks.
@@
F.6 Analysis of the Predictor-Generator-Controller Framework
In our framework, the Time Predictor serves as a crucial component that provides heuristic guidance on system reachability during progress allocation. By estimating the time required for transitions between states, it allows the allocation process to consider not only the logical satisfaction of STL constraints but also the underlying dynamical feasibility between consecutive waypoints, thereby creating favorable conditions for subsequent trajectory generation. The Time Predictor operates in close collaboration with the diffusion-based trajectory generator. Specifically, the predictor estimates the expected travel time between two states, and the generator then produces a trajectory of the corresponding length. Although the true time-to-reach can vary considerably due to stochasticity and unmodeled dynamics, a generator trained on trajectory segments of diverse lengths demonstrates strong generalization. In practice, even a moderately accurate statistical estimate from the predictor typically suffices to produce a dynamically feasible trajectory. Moreover, the feedback control layer further compensates for residual prediction errors, ensuring consistent execution despite model approximation. To empirically validate the effectiveness of this collaboration, we conducted an additional experiment in both Maze2D and AntMaze environments.
Experimental Settings. For each environment, we randomly sampled 1000 start-goal pairs (with goal region radii ranging from 3% to 6% of the arena size). For each case, the Time Predictor estimated the required trajectory length from the start position to the center of the goal region; the trajectory generator then produced a trajectory of that length, which was subsequently executed using a PD controller (in Maze2D) or an inverse dynamics model (in AntMaze) under the strict time-synchronous control protocol described in Section 3.5 (with $k=1$ in Maze2D and $k=2$ in AntMaze). All modules were employed exactly as implemented in our main framework, without additional tuning.
Results and Analysis. The resulting execution success rates, summarized in Table F.5, demonstrate that even in the more challenging AntMaze setting, the current predictor-generator-controller pipeline achieves high overall reliability. These results underscore the effectiveness of modular integration among prediction, generation, and control within our framework. Given this modular design, each component can be further improved independently-for example, by incorporating more accurate time prediction models [51, 52, 58] or more expressive trajectory generators-to handle increasingly complex dynamical systems. We leave such extensions as promising directions for future work.
Table F.5: Execution Success Rate (%) of the Predictor-Generator-Controller Pipeline in Different Environments. Each result is computed over 1000 randomly sampled start-goal pairs.
Environment Umaze Medium Large AntMaze
Execution Success Rate (%) 93.9 89.0 83.7 84.8
F.7 Implementation Details of Experiments
F.7.1 Calculation of the Robustness Value
In our experiments, we compute the robustness values of execution trajectories using the open-source library stlpy [2], which implements quantitative semantics for Signal Temporal Logic (STL). However, in environments such as Maze2D, AntMaze, and Cube, transitions between states often require relatively long trajectories. As a result, the corresponding STL task intervals become lengthy, leading to a substantial computational burden when evaluating robustness directly on the full-resolution trajectories—often exceeding the available computational resources.
To mitigate this issue, we introduce a temporal sampling factor $\eta$, which defines the mapping between the time scale of the STL task and the resolution of the system trajectory used for evaluation. Specifically, one discrete time step in the STL task corresponds to $\eta$ time steps in the system trajectory. When computing robustness, we sample one state every $\eta$ steps to obtain a down-sampled trajectory, and then evaluate robustness on this sampled sequence. Importantly, the STL formula used for evaluation is also temporally rescaled, with all time intervals divided by $\eta$, so that the robustness value is computed with respect to a temporally consistent but shorter-horizon STL specification. This procedure effectively reduces computational overhead while preserving the temporal and logical structure of the original task.
[…]
F.7.2 Experimental Parameter Settings
@@
Some of the parameters involved in the experiments are listed below, and their specific values are shown in Table F.6:
Table F.6: Parameters Used in the Experiments
Env $N_{\mathrm{max}}$ $H$ NN $\gamma$ $\eta$ $k$
Maze2D-Umaze 1 128 64 0.8 8 1
Maze2D-Medium 1 256 256 0.9 8 1
Maze2D-Large 1 384 256 1.1 8 1
AntMaze 1 512 512 1.0 12 2
Cube 1 128 64 1.2 4 1
Custom-Built 1 16&32 64 1 1 1